# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

### v3

@@ I Ternary Classification of Hazard Levels
Table 14: Performance of FOCUS under ternary classification (0: low, 1: medium, 2: high). Results averaged over 3 random seeds.
@@
J Statistical Significance of Performance Gains
@@
For each year (2008, 2019, 2022), we collected the F1-scores from multiple runs (5 seeds per year) and compared the FOCUS results against focal-only using a two-sided Wilcoxon signed-rank test. As shown in Table 15, FOCUS achieved statistically significant improvements in all three years ($p=0.03125$), despite the relatively small number of seeds. Notably, for every random seed, the F1-score achieved with FOCUS was consistently higher than the focal-only counterpart.
@@
Table 15: Wilcoxon signed-rank test comparing F1-scores of FOCUS and focal-only across 5 random seeds for each year.
K Distance-to-Industry Rasters
To capture potential industrial sources of PFAS contamination, we constructed distance rasters for each relevant facility type using data from the EPA’s ECHO database as mentioned in Section 3.1. For each year (e.g., 2008, 2019, 2022), we identified all facilities that were active on or before that year, ensuring historical context was preserved. Facility types were grouped by industrial category (e.g., airports, landfills, chemical manufacturers), and a separate Euclidean distance raster was generated for each category, where pixel values represent the distance to the nearest facility of that type.
@@
L Ablation Results for Image Resolution
Table 16 presents the results of our image resolution ablation study, as highlighted in Section 4.3. We compare the model’s performance when trained on $256\times 256$ versus $512\times 512$ input resolutions, without the use of noise-aware training.
@@
Accuracy 79 $\pm$ 2 75 $\pm$ 3 89 $\pm$ 1 85 $\pm$ 2 84 $\pm$ 3 78 $\pm$ 3
IoU 56 $\pm$ 3 52 $\pm$ 2 67 $\pm$ 2 66 $\pm$ 2 65 $\pm$ 3 56 $\pm$ 2
F-score 70 $\pm$ 3 67 $\pm$ 3 79 $\pm$ 1 78 $\pm$ 2 77 $\pm$ 3 71 $\pm$ 3
Precision 68 $\pm$ 2 66 $\pm$ 3 78 $\pm$ 2 76 $\pm$ 2 74 $\pm$ 3 70 $\pm$ 3
Recall 74 $\pm$ 2 72 $\pm$ 2 80 $\pm$ 1 83 $\pm$ 2 91 $\pm$ 1 84 $\pm$ 2
Table 16: Comparison of performance between $256\times 256$ and $512\times 512$ image resolutions without noise-aware training. Bold values indicate the better result.
M Hyperparameter Tuning and Selection
We tuned the learning rate and batch size empirically by monitoring training stability as outlined in Section 3.3. Specifically, we experimented with learning rates in $\{1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}\}$ and batch sizes in $\{2,4,8\}$. We selected a batch size of 4 and a learning rate of $5\times 10^{-4}$ as they provided the best stability while preserving the benefits of our pretrained backbone, minimizing the risk of large, destabilizing updates. Notably, these final hyperparameter values are consistent with those adopted in the original Prithvi paper, further supporting their suitability in this context.
N Technical Details on Sample Analysis and Training Data for Real-World Predictions
As introduced in Section 4.2, this section provides additional technical details on the real-world PFAS samples collected and analyzed by our team.

### v4

@@ Appendix I Ternary Classification of Hazard Levels
Table 14. Performance of FOCUS under ternary classification (0: low, 1: medium, 2: high). Results averaged over 3 random seeds.
@@
Appendix J Statistical Significance of Performance Gains
@@
For each year (2008, 2019, 2022), we collected the F-scores from multiple runs (5 seeds per year) and compared the FOCUS results against focal-only using a two-sided Wilcoxon signed-rank test. As shown in Table 15, FOCUS achieved statistically significant improvements in all three years ($p=0.03125$), despite the relatively small number of seeds. Notably, for every random seed, the F-score achieved with FOCUS was consistently higher than the focal-only counterpart.
@@
Table 15. Wilcoxon signed-rank test comparing F1-scores of FOCUS and focal-only across 5 random seeds for each year.
@@
Appendix K Distance-to-Industry Rasters
To capture potential industrial sources of PFAS contamination, we constructed distance rasters for each relevant facility type using data from the EPA’s ECHO database. For each year (e.g., 2008, 2019, 2022), we identified all facilities that were active on or before that year, ensuring historical context was preserved. Facility types were grouped by industrial category (e.g., airports, landfills, chemical manufacturers), and a separate Euclidean distance raster was generated for each category, where pixel values represent the distance to the nearest facility of that type.
@@
Figure 12. Example raster channels in an image: (left) land cover, (center) distances from chemical manufacturing industries, and (right) flow direction.
Appendix L Ablation Results for Image Resolution
Table 16 presents the results of our image resolution ablation study. We compare the model’s performance when trained on $256\times 256$ versus $512\times 512$ input resolutions, without the use of noise-aware training.
Table 16. Comparison of performance between $256\times 256$ and $512\times 512$ image resolutions without noise-aware training. Bold values indicate the better result.
@@
Accuracy 37 $\pm$ 2 35 $\pm$ 3 62 $\pm$ 1 59 $\pm$ 2 55 $\pm$ 3 49 $\pm$ 3
IoU 22 $\pm$ 3 20 $\pm$ 2 41 $\pm$ 2 40 $\pm$ 2 36 $\pm$ 3 32 $\pm$ 2
F-score 36 $\pm$ 3 32 $\pm$ 3 57 $\pm$ 1 54 $\pm$ 2 53 $\pm$ 3 50 $\pm$ 3
Precision 54 $\pm$ 2 50 $\pm$ 3 63 $\pm$ 2 57 $\pm$ 2 63 $\pm$ 3 61 $\pm$ 3
Recall 55 $\pm$ 2 51 $\pm$ 2 78 $\pm$ 1 70 $\pm$ 2 74 $\pm$ 1 70 $\pm$ 2
Appendix M Hyperparameter Tuning and Selection
We tuned the learning rate and batch size empirically by monitoring training stability. Specifically, we experimented with learning rates in $\{1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}\}$ and batch sizes in $\{2,4,8\}$. We selected a batch size of 4 because it provided the best stability while preserving the benefits of our pretrained backbone, minimizing the risk of large, destabilizing updates. Notably, these final hyperparameter values are consistent with those adopted in the original Prithvi paper, further supporting their suitability in this context.
Appendix N Technical Details on Sample Analysis and Training Data for Real-World Predictions
This section provides additional technical details on the real-world PFAS samples collected and analyzed by our team.
@@
Appendix O FOCUS Loss Derivation and Assumptions
Setup (latent asymmetric flip noise). Let $z_{i}\in\{0,1\}$ be a latent clean label and $y_{i}\in\{0,1\}$ the observed noisy label. Assume per-pixel flip noise with rate $\eta_{i}\in[0,\tfrac{1}{2})$ (can be asymmetric and spatially varying):
$\Pr(y_{i}=z_{i}\mid x_{i})=1-\eta_{i},\qquad\Pr(y_{i}\neq z_{i}\mid x_{i})=\eta_{i}.$
Let $\pi_{\theta}(x_{i})=\Pr_{\theta}(z_{i}=1\mid x_{i})$ and define $\pi_{\theta}^{(y_{i})}(x_{i})=\pi_{\theta}(x_{i})$ if $y_{i}=1$ and $1-\pi_{\theta}(x_{i})$ if $y_{i}=0$. Then the noisy likelihood is
(7) $\Pr_{\theta}(y_{i}\mid x_{i})=\eta_{i}+(1-2\eta_{i})\,\pi_{\theta}^{(y_{i})}(x_{i}),$
and the per-example noisy negative log-likelihood (NLL) is
(8) $\ell_{\mathrm{noisy}}(\theta;x_{i},y_{i})=-\log\!\Big(\eta_{i}+(1-2\eta_{i})\,\pi_{\theta}^{(y_{i})}(x_{i})\Big).$
Lemma O.1 (Local weighted-CE upper bound).
Fix $\eta_{i}\in[0,\tfrac{1}{2})$ and an anchor $p_{0i}\in(0,1)$. Let $g_{i}(p)=-\log(\eta_{i}+(1-2\eta_{i})p)$. Then for all $p$ in a neighborhood of $p_{0i}$,
(9) $g_{i}(p)\;\leq\;g_{i}(p_{0i})+g_{i}^{\prime}(p_{0i})(p-p_{0i})\;=\;-w_{i}\log p+C_{i},$
where
(10) $w_{i}=\frac{(1-2\eta_{i})p_{0i}}{\eta_{i}+(1-2\eta_{i})p_{0i}}\in[0,1],\qquad C_{i}=g_{i}(p_{0i})+w_{i}\log p_{0i}.$
Interpretation. Lemma O.1 implies the noisy NLL behaves locally like a weighted cross-entropy term $w_{i}\log p$, and the weight $w_{i}$ decreases as $\eta_{i}$ increases (labels become less informative).
Assumption on $M_{i}$ (role of the environmental priors-informed confidence). We use $M_{i}\in[0,1]$ as a practical surrogate for the unknown reliability weight. Concretely, we assume $M_{i}$ is aligned with label cleanliness: (Alignment assumption). We assume that $\mathbb{E}[M_{i}\mid\eta_{i}]$ is non-increasing in $\eta_{i}$, and that $M_{i}$ approximates $w_{i}$ up to a monotone re-scaling:
(11) $\mathbb{E}[M_{i}\mid\eta_{i}]\downarrow\text{ in }\eta_{i},\qquad M_{i}\approx w_{i}.$
Intuitively, $M_{i}$ is larger near locations where supervision is more trustworthy (e.g., near sampling points or along plausible hydrologic pathways) and smaller in regions where point-to-pixel propagation is likely wrong.
We next show that multiplying by a bounded focal factor preserves a local surrogate form.
Lemma O.2 (Focal modulation preserves local surrogate form).
Fix $\gamma\geq 0$ and anchor $p_{0i}\in(0,1)$. There exist constants $A_{i}>0$ and $B_{i}$ such that for all $p$ near $p_{0i}$,
(12) $\log(\eta_{i}+(1-2\eta_{i})p)\;\leq\;A_{i}\,(1-p)^{\gamma}(-\log p)+B_{i}.$
Interpretation. Near a fixed operating point $p_{0i}$, $(1-p)^{\gamma}$ is a positive bounded scalar, so focal loss is a locally rescaled cross-entropy and retains classification-calibrated behavior.
We now connect the above lemmas to the proposed objective.
Theorem O.3 (FOCUS as a noisy-NLL surrogate (local)).
Under the model in (7)–(8) and $\gamma\geq 0$, for each example $i$ there exist constants $\tilde{A}_{i}>0,\tilde{B}_{i}$ such that for all $p$ near $p_{0i}$,
$\ell_{\mathrm{noisy}}(\theta;x_{i},y_{i})\;\leq\;\tilde{A}_{i}\,M_{i}\,(1-p)^{\gamma}(-\log p)+\tilde{B}_{i},\quad\text{where }p=\pi_{\theta}^{(y_{i})}(x_{i}).$
Equivalently, minimizing
(13) $\mathcal{L}_{\text{{FOCUS}}}(x_{i},y_{i})=M_{i}\,(1-\pi_{\theta}^{(y_{i})}(x_{i}))^{\gamma}\,\big(-\log\pi_{\theta}^{(y_{i})}(x_{i})\big)$
minimizes a valid local upper bound on $\ell_{\mathrm{noisy}}$ (up to constants), and thus maximizes a corresponding local lower bound on the noisy log-likelihood.
Connection to expected noisy risk and calibration. Aggregating Theorem O.3 over pixels yields a surrogate for the expected noisy risk $\mathbb{E}[\ell_{\mathrm{noisy}}]$. Since the modulating factor is non-negative and bounded for $p\in(0,1)$ and $\gamma\geq 0$, the objective remains classification-calibrated in the standard sense (locally equivalent to CE).
Edge cases and sanity checks.
  • Clean-label limit. If $\eta_{i}\to 0$, then $w_{i}\to 1$ in (10). If additionally $M_{i}\to 1$, then FOCUS reduces to standard focal loss; with $\gamma=0$ it reduces to CE.
  • Uninformative-label limit. If $\eta_{i}\to\tfrac{1}{2}$, then $w_{i}\to 0$. Under the alignment assumption (11), $M_{i}$ also becomes small, so FOCUS naturally down-weights (or effectively ignores) pixels whose supervision is uninformative.
  • Calibration of $M_{i}$. In practice $M_{i}$ need not equal the true $w_{i}$; it suffices that $M_{i}$ ranks pixels by reliability (monotone alignment). This is precisely the role of our environmental priors-informed construction.
Appendix P Region-Level Risk Signals and Bioaccumulation Evidence
To assess the scientific relevance of model predictions beyond pixel-level accuracy, we aggregate surface-water contamination predictions into region-level risk categories and examine their spatial correspondence with independent fish advisory evidence (Michigan Department of Health and Human Services, nd) over Michigan (Fig 13). Predicted surface-water contamination categories show spatial agreement with regions exhibiting elevated fish contamination, supporting a link between modeled water contamination risk and bioaccumulation-related risk signals.
At the same time, discrepancies between predicted high-risk waters and advisories highlight regions for targeted sampling and scientific follow-up. This analysis illustrates how noise-aware geospatial learning supports AI for Science by integrating large-scale data with sparse observations to guide future measurement efforts.
Figure 13. Region-level aggregation of predicted surface-water PFAS risk over Michigan. Colors indicate low (green), and high (red) predicted risk. Overlaid regions (polylines) correspond to independent fish consumption advisory areas from state monitoring programs.
Region-level visualization comparing aggregated surface-water contamination risk with fish advisory evidence.